# Graph Neural Network Surrogates to leverage Mechanistic Expert Knowledge towards Reliable and Immediate Pandemic Response

### v2

@@ Discussion
In this section, we will introduce the error metric, which we employ to evaluate our surrogate models, discuss the generation of the input data, the mechanistic model which simulates the corresponding reference data, which is used for the model training, as well as the specific architectures of our neural network models.
Performance evaluation and objective function
The output of all models is interpolated to full simulation days so that corresponding demographically and spatially resolved data sets have the same dimension. To evaluate the prediction performance of all neural network-based surrogate models as well as those of the dummy estimator (see next section), we consider the MAPE, which is defined as
as objective function. Here, $n$ is the number of samples, $y_{i}$ are the ground truth daily values and $\widehat{y}_{i}$ are the predicted daily values.
Generation of input data
In order to develop a generalizable surrogate model for a large variety of observed disease dynamics, we need heterogeneous data sets with a large variance in the outcomes, which transfer to the input features of our surrogate models, that eventually cover the full range of all relevant scenarios. The heterogeneity of the generated data sets were assessed qualitatively by visualization and quantitatively by investigating the predictive power of a simple dummy estimator. We compute this dummy estimator by the averaging over 80 % of the data and evaluating its predictive performance on the trajectories of the remaining 20 % of the data using the MAPE given in Eq. (1) as the performance metric. We conclude that the higher the MAPE of the dummy estimator, the higher the variance of the input data.
We use data sets of 10 000 samples for the spatially non-resolved networks and of 1 000 samples for the spatially resolved networks. Given the different scales of population-level data in infectious disease dynamics, in the sense that the number of susceptibles or recovered is generally several magnitudes larger than the number of hospitalized patients or patients in intensive care, we used a logarithmic scaling of the different disease state trajectories.
In order to generate a set of initial conditions and spreading patterns for the surrogate model, we varied the initial segmentation into the infection states as well as the number and intensity of the contact changes. Each individual simulation represented one sample in the data set. The resulting temporal data was split into five days, which is given to the model as input data, and a number of remaining output days data. Each data sample of the GNN is a graph with 400 nodes, each of which has attached the population along all disease states and for all output days in one region; it corresponds to the result of a single simulation run. In other spatio-temporal GNN-based approaches  (?, ?, ?), each graph represents one moment in time and the goal was to predict the graph and its labels in the next time step.
Mechanistic model
@@
The underlying mechanistic expert model which generates the reference training has to integrate the most important disease states and properties of disease dynamics for the surrogate model to learn from. This is required for obtaining a surrogate model that is reliable in practice. While deaths are usually monitored, several indicators such as the number of reported cases, the number of hospitalizations or intensive care unit (ICU) admissions have been considered in Germany to implement or lift nonpharmaceutical interventions against SARS-CoV-2  (?). Consequently, we have chosen an epidemic model based on ordinary differential equations (ODEs) that has already been validated in  (?) and which categorizes subpopulations based on the infection states susceptible, exposed (not infectious), a- or presymptomatic and infectious, mildly symptomatic and infectious, severely infected (hospitalized), critically infected (ICU), dead, and recovered; see Fig. 6. Aside the mentioned indicators, the model particularly considers pre- and asymptomatic transmission as studies on SARS-CoV-2 observed that viral load is peaking close to symptom onset  (?, ?). The model is efficiently implemented in C++ in the MEmilio  (?, ?) framework, which means that it yields very fast model evaluations on a single core. We furthermore divided the population into $G=6$ age groups, that is, from 0 to 4, 5 to 14, 15 to 34, 35 to 59, 60 to 79, and 80+ years.
@@
The transmission process is modeled by the susceptible population $S_{i}(t)$ becoming exposed ($E_{i}(t)$), through infection of either infectious nonsymptomatic ($I_{NS,i}(t)$) or symptomatic individuals ($I_{Sy,i}(t))$, $i\in\{1,\dots,6\}$. This can be described by the ODE
@@
where $\xi_{I_{NS}}$ and $\xi_{I_{Sy}}$ represent the shares of nonisolated a- or presymptomatic and symptomatic infectious individuals, respectively. Furthermore, $\rho_{i}$ denotes the age-resolved probability of becoming infected from a contact with an infectious person, and $\phi_{i,j}(t)$ represents the mean daily number of contacts between individuals of age group $i$ and age group $j$. While, theoretically, $\phi_{i,j}(t)$ could vary arbitrarily, in our realization, it only varies in the context of a contact change point and is constant otherwise; cf. Eq. 3. Additionally, we denote the initial population in age group $j$ by $N_{j}$ and the number of disease-related deaths in age group $j$ up to time $t$ by $D_{j}(t)$, $j=1,\ldots,6$.
The course of disease follows a chain of compartments, where an individual either recovers or moves to an aggravated state. We define $z_{j,i}$ as the $j$-th disease state for age group $i$ and denote the probability of transition from state $z_{j,i}$ to aggravated state $z_{j+1,i}$ by $\mu_{z_{j,i}}^{z_{j+1,i}}$, while consequently $1-\mu_{z_{j,i}}^{z_{j+1,i}}$ is the probability of recovery. The average time an individual spends in a given state $z_{j,i}$ is represented by $T_{z_{j,i}}$. For disease states where progression or recovery is possible, i.e., excluding susceptible, exposed (not infectious), recovered, and dead, transition follows the form
@@
where $z_{j-1,i}$ is the former, less severe state. For the full set of equations, we refer to  (?).
In this study, we use parameters that were obtained for the first waves of wild-type SARS-CoV-2 in 2020. Identical to  (?), the parameters are either directly obtained or calculated based on  (?, Table 2). For the average values over all age groups, the age-specific values are weighted with respect to share of the corresponding age group in the total population provided for Germany by  (?). The overview on the particular values can be found in Tables 1 and 2.
@@
Table 2: Age-resolved parameters for wild-type SARS-CoV-2 based on  (?).
Contact pattern changes are modeled within the the contact rate matrix $\phi(t)=(\phi_{i,j}(t))_{i,j}$. Without loss of generality, let us assume that the first contact change takes place at $c_{1}\in\mathbb{R}$ with $c_{1}>c_{0}$, where $c_{0}$ is the initial time. For a homogeneous reduction factor $0\leq r\leq 1$, we define
@@
where $\phi_{i,j,0}$ is the initial contact rate, $\delta$ defines a transition interval, and $\widehat{\phi}_{i,j}$ represents a transition contact rate to ensure smooth transition, i.e., ${\phi}_{i,j}(t)\in\mathcal{C}^{1}((c_{1}-\varepsilon,c_{1}+\delta+\varepsilon))$ for ${i,j}\in\{1,\ldots,6\}$ with $0<\varepsilon\ll 1$. As suggested in  (?), $\widehat{\phi}$ is realized through a cosine function.
@@
In order to realize the spatial resolution, we employed a generalized graph as described in  (?). Proceeding with this approach, we obtained a graph with one node for each spatial region, which is then equipped with a a local disease dynamics model based on ODEs. Furthermore, we obtained one edge for each pair of age group and disease state, which then implements the corresponding mobility exchange. The 400x400 mobility data for all of Germany’s counties was extracted from a combination of data from the German Federal Employment Agency  (?) and Twitter. The adjacency matrix $A$ needed for the GNN was obtained from this mobility data matrix by setting all nonzero connections to one. While also this metapopulation model was implemented efficiently and did undergo several optimization steps, for large experiments we still require relevant computing resources; c.f. Fig. 5, panel C.
Neural network surrogate model
@@
We conduct a multi-stage model building and optimization process to build up a neural network with spatially resolved infection predictions. Before deriving suitable input data structures for learning a spatially resolved graph-based model with a GNN, we consider the temporal evolution of disease dynamics. For this, we train an MLP, an LSTM and a CNN model, each implemented using the Keras  (?) library with TensorFlow  (?) as backend.
@@
With respect to the temporal evalution, we take several substeps in our tests. In particular, we first consider a population without stratification into age groups. Then, we stratify the population into age groups, for which we consider age-dependent contact patterns and disease parameters. For each model architecture, we perform a grid search on zero to four hidden layers with 32 to 1024 neurons per layer, ranging up to seven layers and 2048 neurons for the final GNN layer type. After selecting a suitable neural network architecture, we fine-tune additional architecture-specific hyper parameters.
@@
After our preliminary tests with simple neural network models which do not consider spectial effects, we turn to GNNs. We use the Spektral library  (?) for defining the graph data and building the GNNs and Keras to train and evaluate the model. Spektral is an open-source Python library based on the Keras API  (?) and TensorFlow 2  (?) implemting GNN models. Other existing libraries, which are not considered here, are PyTorch Geometric  (?) and DeepGraphLibrary  (?).
@@
GNNs can be categorized into spectral and spatial approaches, where spectral approaches have a mathematical foundation in graph signal processing while spatial approaches directly aggregate information from the nodes’ spatial neighborhoods  (?), which is given by the graph. We compare the performance of four different types of GNN layers available in Spektral to build our model. As a first approach, we consider GCNConv layers, which are based on the graph convolutional network (GCN) as proposed by  (?). In this spectral approach, a modified adjacency matrix with self-loops is used. The second type are ARMAConv layers, which are based on  (?). Like the GCNConv layer, this is a spectral approach, and it applies convolutional operations in combination with an autoregressive moving average (ARMA) approach for improved training performance; cf. the discussion below. The third approach uses approximative personalized propagation of neural predictions convolutional (APPNPConv) layers based on  (?), which can also be categorized as a spectral approach with convolutional operations. Lastly, we consider the graph attention convolutional (GATConv) layer from  (?), which is different from the other considered approaches, in the sense that it uses a spatial approach that implements an attention mechanism.
@@
Since the ARMAConv layer performed best (see the Results section), we selected this approach for our surrogate model. Therefore, we briefly provide some details on this specific type of GNN layer as well as our final architecture. In  (?), the authors substituted polynomial filters in GNNs by filters which are based on the ARMA approach. In particular, in our optimized architecture, the final model consists of seven ARMAConv layers with 512 channels per layer and consisting of one stack of GCS layer, which serves as a filter on the node feature input data. Generally, we denote a ARMAConv layer with $k$ GCS stacks as $\textit{ARMA}_{k}$. The $i$th GCS layer in the stack is parametrized by the trainable weight matrix $\textbf{W}_{i}$ as well as the trainable skip connection $\textbf{V}_{i}$, which arises from the the ARMA scheme applied to the GCS stack. one graph convolutional layer with the trainable skip connection $\textbf{V}_{1}$. In the case of the $\textit{ARMA}_{1}$ layer with one GCS stack, the graph convolution layer corresponds to

### v3

@@ 2 Material and methods
In this section, we detail the mechanistic simulator that provides the training targets, introduce the evaluation metric, describe the generation of input data, and document the neural network architectures used in the surrogate models.
2.1 Mechanistic model
@@
The underlying mechanistic expert model which generates the reference training has to integrate the most important disease states and properties of disease dynamics for the surrogate model to learn from. This is required for obtaining a surrogate model that is reliable in practice. While deaths are usually monitored, several indicators such as the number of reported cases, the number of hospitalizations or intensive care unit (ICU) admissions have been considered in Germany to implement or lift nonpharmaceutical interventions against SARS-CoV-2 [35]. Consequently, we have chosen an epidemic model based on ordinary differential equations (ODEs) that has already been validated in [27] and which categorizes subpopulations based on the infection states susceptible, exposed (not infectious), a- or presymptomatic and infectious, mildly symptomatic and infectious, severely infected (hospitalized), critically infected (ICU), dead, and recovered; see Fig. 1. Aside the mentioned indicators, the model particularly considers pre- and asymptomatic transmission as studies on SARS-CoV-2 observed that viral load is peaking close to symptom onset [36, 37]. The model is efficiently implemented in C++ in the MEmilio [38, 39] framework, which means that it yields very fast model evaluations on a single core. We furthermore divided the population into $G=6$ age groups, that is, from 0 to 4, 5 to 14, 15 to 34, 35 to 59, 60 to 79, and 80+ years.
@@
Transmission is modeled via the susceptible population $S_{i}(t)$ becoming exposed ($E_{i}(t)$), through infection of either infectious nonsymptomatic ($I_{NS,i}(t)$) or symptomatic individuals ($I_{Sy,i}(t))$, $i\in\{1,\dots,6\}$. This can be described by the ODE
@@
where $\xi_{I_{NS}}$ and $\xi_{I_{Sy}}$ denote the proportion of nonisolated a- or presymptomatic and symptomatic infectious individuals, respectively. Furthermore, $\rho_{i}$ denotes the age-resolved probability of becoming infected from a contact with an infectious person, and $\phi_{i,j}(t)$ represents the mean daily number of contacts between individuals of age group $i$ and age group $j$. While, theoretically, $\phi_{i,j}(t)$ could vary arbitrarily, in our realization, it only varies in the context of a contact change point and is constant otherwise; cf. Eq. 2. Additionally, we denote the initial population in age group $j$ by $N_{j}$ and the number of disease-related deaths in age group $j$ up to time $t$ by $D_{j}(t)$, $j=1,\ldots,6$.
Disease progression follows a chain of compartments, where an individual either recovers or moves to an aggravated state. We define $z_{j,i}$ as the $j$-th disease state for age group $i$ and denote the probability of transition from state $z_{j,i}$ to aggravated state $z_{j+1,i}$ by $\mu_{z_{j,i}}^{z_{j+1,i}}$, while consequently $1-\mu_{z_{j,i}}^{z_{j+1,i}}$ is the probability of recovery. The average time an individual spends in a given state $z_{j,i}$ is represented by $T_{z_{j,i}}$. For disease states where progression or recovery is possible, i.e., excluding susceptible, exposed (not infectious), recovered, and dead, transition follows the form
@@
where $z_{j-1,i}$ is the former, less severe state. For the full set of equations, we refer to [27].
In this study, we use parameters that were obtained for the first waves of wild-type SARS-CoV-2 in 2020. Identical to [6], the parameters are either directly obtained or calculated based on [27, Table 2]. For the average values over all age groups, the age-specific values are weighted with respect to share of the corresponding age group in the total population provided for Germany by [40]. The overview on the particular values can be found in Tables 1 and 2.
@@
Table 2: Age-resolved parameters for wild-type SARS-CoV-2 based on [27].
Contact pattern changes are modeled within the contact rate matrix $\phi(t)=(\phi_{i,j}(t))_{i,j}$. Without loss of generality, let us assume that the first contact change takes place at $c_{1}\in\mathbb{R}$ with $c_{1}>c_{0}$, where $c_{0}$ is the initial time. For a homogeneous reduction factor $0\leq r\leq 1$, we define
@@
where $\phi_{i,j,0}$ is the initial contact rate, $\delta$ defines a transition interval, and $\widehat{\phi}_{i,j}$ represents a transition contact rate to ensure smooth transition, i.e., ${\phi}_{i,j}(t)\in\mathcal{C}^{1}((c_{1}-\varepsilon,c_{1}+\delta+\varepsilon))$ for ${i,j}\in\{1,\ldots,6\}$ with $0<\varepsilon\ll 1$. As suggested in [27], $\widehat{\phi}$ is realized through a cosine function.
@@
In order to realize the spatial resolution, we employed a generalized graph as described in [27]. Proceeding with this approach, we obtained a graph with one node for each spatial region, which is then equipped with a a local disease dynamics model based on ODEs. Furthermore, we obtained one edge for each pair of age group and disease state, which then implements the corresponding mobility exchange. The $400\times 400$ mobility matrix for all of Germany’s counties was extrapolated from the German Federal Employment Agency [41]. The adjacency matrix $A$ needed for the GNN was obtained from this mobility data matrix by setting all nonzero connections to one. While also this metapopulation model was implemented efficiently and did undergo several optimization steps, for large experiments we still require relevant computing resources; cf. Fig. 6, Panel C.
2.2 Performance evaluation and objective function
From our application, we obtain time-dependent (and spatially resolved) outputs. The output vectors from all models are interpolated to full simulation days so that corresponding demographically and spatially resolved data sets have the same dimension. To evaluate the prediction performance of the neural network-based surrogate models as well as the dummy estimator (see next subsection), we use the mean absolute percentage error (MAPE) as evaluation metric, which is defined as
as objective function. Here, $n$ is the number of samples, $y_{i}$ are the interpolated ”ground truth” daily values and $\widehat{y}_{i}$ are the predicted daily values.
2.3 Generation of input data
To develop a generalizable surrogate model for a large variety of observed disease dynamics, we require heterogeneous data sets with a large variance in the outcomes, which transfer to the input features of our surrogate models and which eventually cover the full range of all relevant scenarios. The heterogeneity of the generated data sets were assessed qualitatively by visualization and quantitatively by investigating the predictive power of a simple dummy estimator. We compute this dummy estimator by the averaging over 80 % of the data and evaluating its predictive performance on the trajectories of the remaining 20 % of the data using the MAPE given in Eq. (3) as the performance metric. We conclude that the higher the MAPE of the dummy estimator, the higher the variance of the input data.
We consider two epidemic regimes that define the initial five input days per sample (values per 100 000 population and stratified by six age groups unless stated otherwise). In the outbreak regime, currently symptomatic individuals are sampled uniformly from $[7,100]$ and recently exposed from $[25,500]$; the remaining compartments are filled consistently to satisfy population balance. In the persistent-threat regime, the share of currently symptomatic individuals is drawn uniformly from $0.01\%,5\%]$; in the worst case, up to $50\%$ of the population may be recently exposed, a-/presymptomatic, or symptomatic. To reflect heterogeneous historical dynamics and immunity, the initial fraction recovered per region is drawn uniformly within the feasible interval implied by the ongoing dynamics so that all compartment shares remain nonnegative and sum to one. The variability of the outbreak and persistent-threat ranges are provided in Panels A and B of Fig. 2. In Panels C) and D) of Fig. 2, we plot the 200 first trajectories for the outbreak scenario (Panel C) and the persistent-threat (Panel D) settings. Here, we see a much more heterogeneous development in Panel D).
Figure 2: Initial conditions over five input days and trajectories over the prediction horizon. A) Initial conditions for outbreak range scenarios. B) Comparison between initial conditions for outbreak range and persistent-threat range scenarios. C) The two hundred first trajectories for the outbreak scenarios. D) The two hundred first trajectories for the persistent-threat scenarios.
Interventions are represented by contact change points. Each sample includes $M\in\{0,1,2,3\}$ changes within the first 30 days; $M$ is drawn uniformly from this set. Change days $(d_{1},\dots,d_{M})$ are drawn without replacement from $\{1,\dots,30\}$ and sorted increasingly, and the homogeneous reduction factors $r_{m}$ are sampled from $\mathcal{U}[0,1)$. Smooth transitions over a short interval $\delta\in(0,1)$ are applied in the mechanistic simulator using a cosine ramp (cf. Eq. 2). Age-structured contact matrices are $6\times 6$ per change point; if $M<3$, the unused slots are masked in the learning input.
We model Germany as a graph with 400 nodes (counties). Each node carries the age-stratified compartment trajectories. The adjacency matrix $A\in\{0,1\}^{400\times 400}$ is derived from the afore-mentioned mobility matrix by mapping nonzero flows to one; this yields a sparse graph with roughly $25\%$ nonzero entries. The same graph is used for training and testing.
Per sample we provide the first five days as input and predict subsequent horizons of 30, 60, or 90 days. For non-spatial baselines, we use a 2D array of dimension $5\times 162$ array with the five input days defining the rows each with $48$ age-resolved compartment values (six age groups times eight compartments) plus up to three $6\times 6$ contact matrices with their three contact change days and reduction factors. For the GNNs, each sample consists of node features $X\in\mathbb{R}^{400\times 354}$ containing the five input days of compartment values ($48$ per node) and the intervention descriptors (contact matrices, change days, reduction factors) broadcast per node, paired with the adjacency matrix $A$. This schema mirrors the input layout illustrated in Fig. 3. To balance scale differences across compartments (e.g., Susceptible vs. ICU), we apply a $\log(1+x)$ transform to all count-valued series before training; evaluation is performed on the original scale after inverse transforming.
Figure 3: Non-spatial vs. spatial input encoding. A) Non-spatial 2D input (5 $\times$ 162): five input days with 48 (age groups $\times$ compartments) values plus up to three $6\times 6$ contact matrices, their change days, and reduction factors. B) Spatial GNN input: for each node (400 counties) the same features as in Panel A are broadcast and paired with the adjacency matrix.
We generate 10,00010{,}000 samples for the non-spatial baselines and 1,0001{,}000 samples for the spatial GNNs. Unless stated otherwise, we use an 80–10–10 split into train/validation/test on the sample level and perform five-fold cross-validation on the training split during architecture search to avoid split artifacts; the reported test MAPE refers to the held-out test set. The dummy estimator is computed on the training split and evaluated on the held-out portion within each regime as a single-number proxy for input heterogeneity; detailed values are reported in the Results.
2.4 Neural network surrogate model
@@
We conduct a multi-stage model building and optimization process to build up a neural network with spatially resolved infection predictions. Before deriving suitable input data structures for learning a spatially resolved graph-based model with a GNN, we consider the temporal evolution of disease dynamics. For this, we train an MLP, an LSTM and a CNN model, each implemented using the Keras [42] library with TensorFlow [43] as backend.
@@
With respect to the temporal evaluation, we take several substeps in our tests. In particular, we first consider a population without stratification into age groups. Then, we stratify the population into age groups, for which we consider age-dependent contact patterns and disease parameters. For each model architecture, we perform a grid search on zero to four hidden layers with 32 to 1024 neurons per layer, ranging up to seven layers and 2048 neurons for the final GNN layer type. After selecting a suitable neural network architecture, we fine-tune additional architecture-specific hyperparameters.
@@
After our preliminary tests with simple neural network models which do not consider spatial effects, we turn to GNNs. We use the Spektral library [44] for defining the graph data and building the GNNs and Keras to train and evaluate the model. Spektral is an open-source Python library based on the Keras API and TensorFlow 2 [43] implementing GNN models. Other existing libraries, which are not considered here, are PyTorch Geometric [45] and DeepGraphLibrary [46].
@@
GNNs can be categorized into spectral and spatial approaches, where spectral approaches have a mathematical foundation in graph signal processing while spatial approaches directly aggregate information from the nodes’ spatial neighborhoods [47], which is given by the graph. We compare the performance of four different types of GNN layers available in Spektral to build our model. As a first approach, we consider GCNConv layers, which are based on the graph convolutional network (GCN) as proposed by [48]. In this spectral approach, a modified adjacency matrix with self-loops is used. The second type are ARMAConv layers, which are based on [49]. Like the GCNConv layer, this is a spectral approach, and it applies convolutional operations in combination with an autoregressive moving average (ARMA) approach for improved training performance; cf. the discussion below. The third approach uses approximative personalized propagation of neural predictions convolutional (APPNPConv) layers based on [50], which can also be categorized as a spectral approach with convolutional operations. Lastly, we consider the graph attention convolutional (GATConv) layer from [51], which is different from the other considered approaches, in the sense that it uses a spatial approach that implements an attention mechanism.
@@
Since the ARMAConv layer performed best (see the Results section), we briefly provide some details on this specific type of GNN layer as well as our final architecture. In [49], the authors substituted polynomial filters in GNNs by filters which are based on the ARMA approach. In particular, in our optimized architecture, the final model consists of seven ARMAConv layers with 512 channels per layer and consisting of one stack of GCS layers, which serves as a filter on the node feature input data. Generally, we denote a ARMAConv layer with $k$ GCS stacks as $\textit{ARMA}_{k}$. The $i$th GCS layer in the stack is parametrized by the trainable weight matrix $\textbf{W}_{i}$ as well as the trainable skip connection $\textbf{V}_{i}$, which arises from the ARMA scheme applied to the GCS stack. In the case of the $\textit{ARMA}_{1}$ layer with one GCS stack, the graph convolution layer corresponds to